OpenStudy (anonymous):

The product of 2 consecutive odd numbers is 63. What is the larger number? @ex0rcist_6_6_6 @noemy_bieber

OpenStudy (cwrw238):

x(x+2) = 63

OpenStudy (cwrw238):

solve this to get the lower number , the larger will be = x+2

OpenStudy (anonymous):

wht do i do

OpenStudy (cwrw238):

right x(x+2) = 63 x^2 + 2x - 63 = 0 can you solve this trinomial?

OpenStudy (anonymous):

no

OpenStudy (cwrw238):

(x + 9(x - 7) = 0 x = -9 or 7 ignore the -9 x = 7 the number you want is 7 + 2 = 9

OpenStudy (cwrw238):

gotta go